I need to find the total cost or sale price to the nearest cent.
I have no idea how to do it.
Here is the question:
$24.50 sandals, 20% discount.

Answer by rfer(16322)

24.50*0.80=$19.60
---------------------
or
24.50*0.20=4.90
24.50-4.90=$19.60